OpenStudy (kaliyah0524):

fan & medal! Graph the linear function by finding​ x- and​ y-intercepts. Then write the equation using function notation. x=4y

OpenStudy (wholock):

Start by putting it in slope-intercept form. y=mx + b Look at it like this: 4y = x

OpenStudy (kaliyah0524):

okay. What next

OpenStudy (wholock):

You get y = 1/4x now make x = 0 to find the y-intercept.

OpenStudy (wholock):

Then set y to 0 to get the x-intercept.